What is the y value of
y = x^2 - 4x
when the value of x is -1 negative 1
or the value of is ix 1 positive 1
are they the same? why or why not?
can you show how you solved it?
Thanks so much.

Answer by rfer(16322)

y=x^2-4x
y=-1^2-4*-1
y=1+4
y=5
-----------
y=1^2-4*1
y=1-4
y=-3